OpenStudy (anonymous):

Solve for the base. 36 is 0.75% of

OpenStudy (ipwnbunnies):

What base? ..... So, we're saying x*0.0075 = 36. Can you solve for x?

OpenStudy (anonymous):

yes solving for x. how come you moved the decimal point?

OpenStudy (ipwnbunnies):

It's just a way of writing a percent when doing a problem. For example, we want to 50% of 10. Would you agree 50% of 10 is 5?

OpenStudy (ipwnbunnies):

want to find*

OpenStudy (anonymous):

well yes 5 is 50%. but that doesn't explain why you moved the decimal point. I just tried it on my homework and 36 isn't right.

OpenStudy (kc_kennylau):

Are you aware that 1%=0.01

OpenStudy (kc_kennylau):

Or 100%=1

OpenStudy (kc_kennylau):

That's why 0.75%=0.0075

OpenStudy (ipwnbunnies):

Stay with me lol. If we wrote it out in a problem, it'll be 10 * 0.5 = 5. For percents, move the decimel 2 spaces to the left.

OpenStudy (kc_kennylau):

So you're asked to find 36 is 0.75% of 'what'

OpenStudy (kc_kennylau):

That means 0.75% of 'what' is 36

OpenStudy (kc_kennylau):

Or 0.75%\(\times\)'what'=36

OpenStudy (kc_kennylau):

Which is the same as 0.0075'what'=36

OpenStudy (kc_kennylau):

Get it?

OpenStudy (anonymous):

yes that makes sense! Thank you for explaining all of that! I really appreciate it

OpenStudy (kc_kennylau):

no problem